Give the reasons for the steps in the following proof. If 3x+(-1)=0, then x= 1/3
3x+(-1)=0
[3x+(-1)]+1= 0+1
3x+[(-1)+1]= 0+1
3x+0= 0+1
3x= 1
1/3(3x)= 1/3(1)
(1/3*3)x= 1/3*1
1*x= 1/3*1
x= 1/3

Answer by Edwin McCravy(20086)


Give the reasons for the steps in the following proof. 

If 3x+(-1)=0, then x = 1/3

3x+(-1)=0                     GIVEN
[3x+(-1)]+1= 0+1              EQUALS ADDED TO EQUALS GIVES EQUALS 
3x+[(-1)+1]= 0+1              ASSOCIATIVE PRINCIPLE OF ADDITION
3x+0= 0+1                     ADDITIVE INVERSE PROPERTY
3x= 1                         ADDITIVE IDENTITY PROPERTY
1/3(3x)= 1/3(1)               EQUALS MULTIPLIED BY EQUALS GIVES EQUALS  
(1/3*3)x= 1/3*1               ASSOCIATIVE PRINCIPLE OF MULTIPLICATION
1*x= 1/3*1                    MULTIPLICATIVE INVERSE PROPERTY  
x= 1/3                        MULTIPLICATIVE IDENTITY PROPERTY

Edwin